OpenStudy (anonymous):

please help, i am all left out here Two dices are tossed together, what is the probability of having a multiple of 12

OpenStudy (anonymous):

6

OpenStudy (anonymous):

that is not right. please show your work

OpenStudy (anonymous):

since u rolled 2 dices 12/2 equals 6!

OpenStudy (anonymous):

@Ilovecake help please

OpenStudy (ilovecake):

Probability for rolling two dice with the six sided dots such as 1, 2, 3, 4, 5 and 6 dots in each die. When two dice are thrown simultaneously, thus number of event can be 62 = 36 because each die has 1 to 6 number on its faces. Then the possible outcomes are shown in the below table. You are right @rocker20026

OpenStudy (ilovecake):

He is right.

OpenStudy (anonymous):

but the options i have are I/3 1/9 2/3 1/6

OpenStudy (ilovecake):

then it is 1/6: http://www.math-only-math.com/images/probability-for-rolling-two-dice.jpg

OpenStudy (anonymous):

can i post more question please? will medal and fan

OpenStudy (anonymous):

What is the probability of having sum of seven in the throwing of a pair of dice once?

OpenStudy (ilovecake):

ok, but can you post it on a new post? thanks

OpenStudy (anonymous):

you there?

OpenStudy (ilovecake):

yes

OpenStudy (ilovecake):

post this on a new question.

OpenStudy (tkhunny):

Ummm... The only multiple of 12 that is possible is 12. P(12) = 1/36

OpenStudy (ilovecake):

I agree with @tkhunny

OpenStudy (anonymous):

so is it 12/36?

OpenStudy (tkhunny):

No, p(12) = 1/36 Where did you get 12/36? There are 36 different outcomes. Only ONE of them is 12.

OpenStudy (anonymous):

but the options given does not contain 1/36

OpenStudy (ilovecake):

no I was saying yea for the other part.

OpenStudy (anonymous):

then what is the answer?

OpenStudy (tkhunny):

So? When you have the right answer, how much does it matter that the answer choices are wrong? p(12) = 1/36 when rolling two fair dice. You'll need another option.

OpenStudy (tkhunny):

You can get 7 various ways. 1 + 6 2 + 5 3 + 4 4 + 3 5 + 2 6 + 1 That's 6, so Pr(7) = 6/36 = 1/6

OpenStudy (anonymous):

ok. what about this Four elements (A,B,C, and D)required to produce a drug for Ebola virus treatment are to be mix in the ratio 1:2:4:5, if this information is to represented on a pie chart, what angle will element C represented with on the pie chart

OpenStudy (anonymous):

am lost in that question

OpenStudy (phi):

say you mix in 1 cup of "A" how many cups of B, C and D do you add in ?

OpenStudy (anonymous):

please help me. dont really understand the steps

OpenStudy (phi):

the ratio 1:2:4:5 tells you how much of A, B, C and D to mix together if you start with 1 cup of A, then you add 2 cups of B, how many cups of C ?

OpenStudy (anonymous):

ok. please continue am learning through your process

OpenStudy (phi):

how many cups of C ?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

am i correct?

OpenStudy (phi):

and 5 cups of D

OpenStudy (phi):

next, how many cups total does all that add up to ?

OpenStudy (anonymous):

12 total

OpenStudy (phi):

next, we know 4 cups of that total is C what fraction of the total is C ?

OpenStudy (anonymous):

4/12

OpenStudy (phi):

which simplifies to 1/3 now to the question. if you make a circle, and divide it into 3 equal sections  any idea what the angle is ?